OpenStudy (anonymous):

The table below shows some values of f(x) and g(x) for different values of x: x f(x) = 3x + 7 g(x) = 4)-x -2 16 -1 4 0 1 1 2 13 Complete the chart and determine the solution of the equation f(x) = g(x) x = -2 x = -1 x = 1 x = 4

OpenStudy (astrophysics):

What's stopping you from plugging in the numbers?

OpenStudy (astrophysics):

Plug in the given x values and evaluate the functions

OpenStudy (anonymous):

some help u r lol

OpenStudy (anonymous):

jdoe help me please how do i determine the solution

OpenStudy (anonymous):

i plugge in the numbers

OpenStudy (jdoe0001):

hmmm well. "complete the chart" is just simply setting "x" to the value on the left column for each function and getting the "y"

OpenStudy (anonymous):

ik i already have the it set

OpenStudy (jdoe0001):

and once you complete the chart you'd notice that, at some point f(x) will give a value that resembles what g(x) gave, over the same row at that row, f(x) = g(x)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

then the answer is x=-1 :3

OpenStudy (anonymous):

i think its talking about the row or the answer its self because f(4)=g(4)

OpenStudy (jdoe0001):

yeap x = -1 f(x) = g(x)

OpenStudy (anonymous):

k thnx

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

help me with one more plz?

OpenStudy (anonymous):

well 2 more lol